# FEUDA: Frustratingly Easy Prompt Based Unsupervised Domain Adaptation

### v1

@@ 5.2 Models and Training
We train each training phase for 30,000 steps on MNLI and 15,000 steps on the Amazon dataset. We train with Adam and use a batch size of 8, learning rate of 0.003. We set the maximum sequence length to 256 tokens. We use length normalization during evaluation, as proposed by Liu et al. (2022). For each experiment, we report the mean and standard deviation across 3 runs. More details can be found in Appendix D.
5.3 Baselines
We compare FEUDA with three baselines below.
  • Src Only: We fine-tune the model on the source labeled data, in a single training phase.
@@ 5.3 Baselines
  • Src+Tgt (All labeled): We fine-tune the model for classification, using labeled data from both the source and target domains. This serves as an upper bound on target domain performance.
@@
  • MMD: The current state of the art for UDA in a PEFT setup promotes domain invariance by maximum mean discrepancy (MMD) Malik et al. (2023). This method measures the multi-kernel maximum mean discrepancy (MK-MMD) Gretton et al. (2012); Bousmalis et al. (2016) between source and target embeddings from each transformer layer and sums them to obtain an aggregate loss $\mathcal{L}_{\text{div}}$. The final loss is the weighted sum of $\mathcal{L}_{\text{div}}$ and the classification loss, i.e., $\mathcal{L}=\lambda\;\mathcal{L}_{\text{cls}}+(1-\lambda)\;\mathcal{L}_{\text{div}}$, where $\lambda$ gradually changes from 0 to 1 during training.
6 Results & Analysis
6.1 FEUDA is Competitive for UDA
We compare our method with other baselines over 24 domain pairs (16 additional pairs in Appendix E) domain pairs across the MNLI and Amazon product review datasets. In these experiments, we use the T5v1.1 Base model and (IA)${}^{3}$ PEFT method.
FEUDA outperforms methods that explicitly promote domain invariance
@@ 6.1 FEUDA is Competitive for UDA
Table 1 shows the classification accuracy on the MNLI dataset. We find that FEUDA is a competitive method to MMD. For example, for Travel (T) $\to$ Government (G), FEUDA yields an accuracy of 83.6% on the target domain, equalling the upper bound of the Src+Tgt baseline. In comparison, MMD yields an accuracy of only 79.3%. Additionally, FEUDA performs more stably than MMD. For example, for Fiction (F) $\to$ Government (G), minimizing MMD yields a variance of over 20% across runs. This observation is consistent with existing findings Kashyap et al. (2021); Han and Eisenstein (2019) that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, result in training instabilities and vanishing gradients. We discuss the performance of the MMD method in more detail in Appendix J.
We also see similar results on the Amazon dataset in Table 2 (full results in Appendix E). For example, for Apparel (A) $\to$ Movies (M), FEUDA yields an accuracy of 83.3% on the target domain, approaching the upper bound of the Src+Tgt baseline. In comparison, MMD yields an accuracy of 81.3%. A visualization of sentence embeddings in Figure 4 (Appendix E) suggests that representations learned through FEUDA are not domain invariant. In Appendix I, we show FEUDA outperforms additional baselines, including DANN Ganin et al. (2016), weight interpolation Ilharco et al. (2022), and domain divergence minimization by Wasserstein distance.
6.2 Analyzing the Impact of MLM on UDA
In this section, we aim to understand how MLM training on the source and target domains boosts classification on the unlabeled target domain.
Impact of Masked Words
We hypothesize that by having to predict certain masked words during MLM training, the model implicitly learns information about the classification task on the unlabeled domain. For example, given the masked sentence,
“I really _ the movie, it was a fascinating watch.”
The only way the model can predict the masked word is by using the sentence context and identifying the sentiment of certain words. In this case, the word “fascinating” implies a positive sentiment, so the model may predict the masked word to be a positive word like “loved” or “enjoyed”. Thus, the model would implicitly learn information about the downstream task, by predicting masked words (e.g., “fascinating”) indicative of the class label.
@@ 6.2 Analyzing the Impact of MLM on UDA
To test this hypothesis, we quantize the “informativeness” of each word to a classification task. An informative word is one that is highly correlated with any of the labels in the downstream task. Specifically, we follow Gururangan et al. (2018) and use pointwise mutual information (PMI) Fano (1961) of the word with respect to the class label:
@@
where we count the frequency of a word-class pair on the labeled data $\mathcal{D}_{\text{src}}$ to estimate $p(\text{word},\text{class})$, and similarly estimate $p(\text{word})$ and $p(\text{class})$ by counting a word and a class individually on $\mathcal{D}_{\text{src}}$. These informative words are similar to pivot features (Blitzer et al., 2006; Ben-David et al., 2020, inter alia), with the exception that they are chosen based on information from the source domain only.
To compare with random masking of $k\%$ words, we selectively mask the top $k\%$ or bottom $k\%$ of informative words in a sentence, ranked by their PMI with any inference label ($k=15$). We also filter out low-frequency words from the selection. [^3] Any word that occurs less than 10 times in the entire training corpus is considered to be low frequency. We use the T5v1.1 base model with (IA)${}^{3}$ on the Apparel $\to$ Movies pair for analyzing the impact of masked words at inference and pre-training time.
Impact at Inference: We use a prompt-based classifier trained by FEUDA to classify three versions of a test set at inference: the original test set and the other two versions with informative and uninformative words masked respectively. Figure 1 (corresponding Table 10 in Appendix G) shows us that the presence of both informative and uninformative words are essential for strong classification performance, with performance being highest on the original unmasked sequences. Interestingly, the source-domain performance is only hurt by the masking of informative words, confirming that these words are highly indicative of the downstream classes.
Impact at Pre-training: To further confirm the phenomenon, we alter the masking strategies in the MLM pre-training phase of FEUDA. We compare the original random masking with informative and uninformative masking, maintaining a fixed masking rate (15%) across masking strategies. Table 3 confirms that random masking is most helpful for target-domain classification. To isolate any effects of PEFT methods or pre-training data, we repeat the analysis by fine-tuning the instruction-tuned Flan-T5 Chung et al. (2022) and notice a similar trend (Table 11 in Appendix G). We hypothesize that the model learns semantic features through the masking of informative words and background features through the masking of uninformative words, and both sets of features are essential for classification on the unlabeled domain.
Masking Strategy Accuracy
Source Target
Random 93.5 (0.4) 83.3 (0.5)
Informative 93.3 (0.5) 78.3 (0.7)
Uninformative 92.9 (0.1) 79.6 (0.6)
Table 3: Impact of word selection for masking during training. Masking words at random is more powerful than selectively masking informative or uninformative words. This indicates that the model requires both semantic features (learnt through masking informative words) and background features (learnt through masking uninformative words) for classification on the unlabelled target domain.
Figure 1: Impact of masking at inference. We evaluate FEUDA on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task. The performance of the model is best with the original unmasked sequences, indicating the presence of both informative and uninformative words are essential.
Impact of Masking Rate
While masking 15% of a sequence is considered standard for random masking, previous work has shown that BERT-sized models Devlin et al. (2019) can learn from as high as 80% masking rates during pre-training followed by adaptation to a labeled task. Here, we explore the role of masking rates during continued pre-training for adaptation to an unlabeled task.
Figure 2 (Table 12 in Appendix) shows the impact of varying the masking rate on the source and target domains. With masking rates under the optimal value of 15%, the semantic and background features learned through model prediction of masked words is limited, hurting performance on the target domain. Beyond the 15% rate, the classification performance on the source domain is largely maintained, even at a 90% masking rate, matching previous findings Wettig et al. (2023). However, the performance on the target domain rapidly decreases with an increasing masking rate.
To explain the performance drop, we hypothesize that since the model never sees any labeled data of the target domain, it heavily depends on the signal it gets from the unlabeled data through masking. Effectively masking a majority of a sequence removes the background and semantic features from the sequence, both of which are necessary for downstream classification on the domain.
Figure 2: Impact of Masking Rate on FEUDA. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
6.3 Extensions to More Settings
Method Source Accuracy Target Accuracy
Src Only Src+Tgt MMD FEUDA Src Only Src+Tgt MMD FEUDA
Fine-Tuning 93.9 (0.5) 94.0 (0.4) 93.8 (0.3) 95.0 (0.4) 82.0 (1.1) 86.4 (0.4) 82.4 (1.6) 84.4 (0.3)
(IA)3 93.7 (0.1) 94.1 (0.3) 94.7 (0.3) 93.5 (0.4) 81.3 (1.4) 85.8 (0.5) 81.3 (0.6) 83.3 (0.5)
Adapters 93.6 (0.1) 94.6 (0.3) 94.4 (0.7) 94.3 (0.2) 80.8 (1.3) 85.3 (0.5) 79.1 (0.3) 82.7 (0.5)
Table 4: Performance of FEUDA across different adaptation methods with the T5v1.1 base model on the Apparel $\to$ Movies domain pair. FEUDA remains more powerful than MMD across all methods.
Prompting is common practice with large language models which are too compute-intensive for traditional fine-tuning. For the same reasons, the setup is also frequently combined with learning from limited examples and parameter-efficient fine-tuning. In this section, we explore these settings, using the Apparel (A) $\to$ Movies (M) domain pair from the Amazon Reviews dataset.
Model Types & Scales
We evaluate the performance of FEUDA over larger and instruction tuned encoder-decoder models, using T5v1.1 XL and T0 Sanh et al. (2022). Table 9 (Appendix F) shows a wider gap between MMD and FEUDA with higher model capacity, and this gap is further increased with instruction tuning. We also explore the utility of FEUDA when using causal language modeling (CLM) with a decoder-only language model. Table 5 shows us that FEUDA provides strong improvements on both domains, equalling the Src+Tgt baseline on the target domain.
Method Accuracy
Source Target
Src Only 86.9 (1.4) 65.9 (1.4)
Src+Tgt 86.8 (1.0) 73.4 (0.1)
MMD 86.9 (0.9) 66.8 (0.8)
FEUDA 89.3 (0.5) 73.5 (0.8)
Table 5: Performance of FEUDA with causal language modeling and decoder-only architectures, with the GPT-2 medium model on the Apparel $\to$ Movies domain pair. FEUDA remains powerful, and improves performance on the source and target domains.
Adaptation Methods
PEFT approaches have been shown to introduce resilience to domain shift Fu et al. (2023). To isolate this effect from the FEUDA framework, we evaluate our method in a full-fine-tuning setup. Further, we compare with two PEFT approaches: Adapters Houlsby et al. (2019) and (IA)${}^{3}$ Liu et al. (2022). We choose Adapters because He et al. (2022) present a unified view of PEFT approaches which shows that the operations applied by Adapters are very similar to those of Prefix Tuning Li and Liang (2021) and LoRA Hu et al. (2022). We choose (IA)${}^{3}$ since it is a state-of-the-art PEFT approach that uses a fraction of the learnable parameters of Adapters. Table 4 shows FEUDA beats MMD across different adaptation methods. We also note that fine-tuning yields slightly better performance on both domains for all UDA methods.
Few-Shot Learning
For the following experiments, we assume access to $k$ labeled source domain examples. For FEUDA, we assume access to the full unlabeled dataset in both domains for MLM pre-training, and $k$-shot access to labeled source data points for the second phase of supervised training. For a fair comparison, we also introduce a two-phase version of the MMD pipeline – the first phase minimizes MMD between unlabeled source and target domain embeddings (full data access), while the second phase optimizes supervised training on the source domain ($k$-shot).
Figure 3 (Table 13 in Appendix H) showcases FEUDA clearly outperforming both variants of MMD, across three different models. In Appendix H, we also repeat the analysis from Section 6.2 and find that, like with the full-data setting, semantic and background features are required for classification on the unlabeled target domain. However in this setting, downstream classification is aided more by the masking of informative words, rather than uninformative words.
Figure 3: Target domain accuracy of FEUDA across different models, in a 256-shot learning setup on the Apparel $\to$ Movies domain pair. We see FEUDA retaining strong performance on the target domain across models.

### v2

@@ 3.1 Experimental Setup
Each training phase is 30,000 steps long for MNLI and 15,000 steps for the Amazon dataset. We use Adam, a batch size of 8, and a learning rate of 0.003. We set the maximum sequence length to 256 tokens. We use length normalization during evaluation, as proposed by Liu et al. (2022). For each experiment, we report the mean and standard deviation across 3 runs. More details can be found in Appendix B.
Baselines
Since our goal is to study the behaviour of CPT for generative UDA, we it compare with a simple supervised baseline, and a state-of-the-art invariance based approach.
@@
  • Src+Tgt (All labeled): We fine-tune the model using labeled data from both the source and target domains. This serves as an upper bound on target domain performance.
@@
  • UDAPTER: Malik et al. (2023) propose an invariance based method that measures the multi-kernel maximum mean discrepancy (MK-MMD) (Gretton et al., 2012; Bousmalis et al., 2016) between source and target embeddings from each transformer layer and sums them to obtain an aggregate loss $\mathcal{L}_{\text{div}}$. The final loss is the weighted sum of $\mathcal{L}_{\text{div}}$ and the classification loss, i.e., $\mathcal{L}=\lambda\;\mathcal{L}_{\text{cls}}+(1-\lambda)\;\mathcal{L}_{\text{div}}$, where $\lambda$ gradually changes from 0 to 1 during training. Here, we use the embeddings from a model as it is being instruction tuned on the downstream classification task. Their method achieves state-of-the-art performance, and outperforms popular UDA approaches like DANN (Ganin et al., 2016) and DSN (Bousmalis et al., 2016); thus we only compare CPT with this approach.
3.2 Continued Pre-training is Competitive with Domain-Invariance Methods
Dataset Src+Tgt UDAPTER CPT
Amazon 92.66 (0.45) 89.02 (2.17) 89.34 (0.48)
MNLI 78.14 (0.25) 70.19 (1.71) 74.12 (0.68)
Table 1: Avg. target-domain classification accuracy and standard deviation over 3 runs on Amazon Review and MNLI.
Performance
We compare CPT with other baselines over 40 domain pairs of the MNLI and Amazon Review datasets, and report the average accuracies over all pairs in Table 1. We see that CPT is competitive to UDAPTER. (Appendix C contains detailed results over 40 pairs and significance tests to check for competitiveness.) Interestingly, a visualization of sentence embeddings in Figure 8 (Appendix C) suggests that representations learned through CPT are not domain invariant. In addition to the MMD based method of Malik et al. (2023), we also compare CPT on one domain pair with other methods that promote domain invariance (DANN (Ganin et al., 2016), CORAL (Sun et al., 2017)) and weight interpolation (Ilharco et al., 2022) in Appendix C, further confirming the competitiveness of CPT.
Stability
@@ 3.2 Continued Pre-training is Competitive with Domain-Invariance Methods
CPT performs more stably than UDAPTER, with the invariance-based method often reporting a variance of over 20% across runs (Table 6 in Appendix C). For example, for the MNLI pair Fiction (F) $\to$ Government (G), minimizing UDAPTER yields a variance of 23.4% across runs. This observation is consistent with existing findings (Kashyap et al., 2021; Han & Eisenstein, 2019) that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, result in training instabilities and vanishing gradients. We discuss this in more detail in Appendix J.
4 How General are the Benefits of CPT?
[…]
CPT Helps Decoder-only Models.
We extend our analysis from MLM with encoder-decoder language models, to causal language modeling (CLM) with decoder-only language models, using GPT-2 (medium) (Radford et al., 2019). We perform CLM in the first training phase by simply training the model for next-word prediction given the original sequence. Table 2 shows that CPT provides strong improvements on the target domain in comparison to the invariance based baseline.
CPT Outweighs Invariance-based Methods for Instruction-tuned Models.
We evaluate the performance of CPT over T5v1.1 XL (3B parameters) and the instruction-tuned T0 (3B parameters) (Sanh et al., 2022). Figure 2 (Table 8 of Appendix D) shows a wider gap between UDAPTER and CPT with higher model capacity, and this gap is further increased with instruction tuning. We hypothesize that this gap is due to the vast difference between the objectives of domain invariance and instruction tuning.
Method Accuracy
Src+Tgt 79.8 (0.3)
UDAPTER 66.0 (1.1)
CPT 75.8 (0.4)
Table 2: Performance of CPT with causal language modeling for the decoder-only GPT-2 model. CPT significantly outperforms the invariance-based method.
Figure 1: The performance gap between CPT and UDAPTER increases with larger models, from T5v1.1 Base (60M parameters) to T5v1.1 XL (3B parameters), and further increases with instruction tuning (T0 3B).
Figure 2: Performance of CPT across different tuning approaches with the T5v1.1 base model. CPT remains more powerful than UDAPTER across all tuning approaches.
CPT Benefits are Consistent across Tuning Approaches.
PEFT approaches have been shown to introduce resilience to domain shift (Fu et al., 2023). To isolate this effect from the CPT framework, we use T5v1.1 to evaluate CPT in a full fine-tuning setup. Additionally, we compare CPT with two PEFT approaches [^4] We choose Adapters because He et al. (2022) present a unified view of PEFT approaches which shows that the operations applied by Adapters are very similar to those of Prefix Tuning (Li & Liang, 2021) and LoRA (Hu et al., 2022). We choose (IA)3 since it is a state-of-the-art PEFT approach that uses a fraction of the learnable parameters of Adapters (More in Appendix E). : Adapters (Houlsby et al., 2019) and (IA)3 (Liu et al., 2022). We see in Figure 2 (Table 9 in Appendix E) that CPT continues to perform stronger than the domain invariance based UDAPTER method.
CPT Outperforms Invariance-based Methods in Low-data Regimes.
In this low-data experiment, we assume access to $k$ labeled source-domain examples. For CPT, we assume access to the full unlabeled dataset in both domains for the first training phase, and $k$-shot access to labeled source-domain examples for the second phase of supervised training. For a fair comparison, we also introduce a two-phase version of the UDAPTER pipeline—the first phase minimizes MMD between unlabeled source and target domain embeddings (full data access), while the second phase optimizes supervised training on the source domain ($k$-shot). Figure 4 (Table 10 in Appendix F) showcases CPT clearly outperforming both variants of UDAPTER, across three different models for $k=256$. Furthermore, Figure 4 (Table 11 in Appendix F) shows CPT providing consistent improvements in as low as 32-shots, unlike the unstable invariance-based approach.
Figure 3: Few-shot performance of CPT, for varying $k$. Unlike both variants of UDAPTER, CPT is stable and provides consistent benefits across number of shots.
Figure 4: Performance of CPT across different models, in a 256-shot learning setup. Unlike both variants of UDAPTER, CPT is stable and provides consistent benefits across models.
5 To What Extent is Target-Domain Exposure Beneficial?
Given the vast distributions language models are pre-trained on, a natural assumption might be that the model has already been exposed to a domain similar to the target domain during pre-training. This would mean that the model could simply extrapolate the learned downstream task from the source to the target domain, questioning the need for CPT.
[…]
Figure 5: Impact of Masking Rate on CPT. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
For a more fine-grained analysis, we investigate the impact of degree of exposure to the target domain, by varying the masking rates during the first phase of training. While masking 15% of a sequence is considered standard for random masking, previous work has shown that BERT-sized models (Devlin et al., 2019) can learn from as high as 80% masking rates during pre-training followed by adaptation to a labeled task (Wettig et al., 2023). The source-domain performance shown in Figure 5 (Table 12 in Appendix) matches this trend. However, high masking rates effectively reduce the exposure of the model to target data, strongly deteriorating the performance on the target domain [^5] With masking rates under the optimal value of 15%, the semantic and background features learned through model prediction of masked words is limited, hurting performance on the target domain.. We hypothesize that since the model never sees any labeled data of the target domain, it heavily depends on the signal it gets from the unlabeled data through masking.
Figure 6: Vocabulary overlap between label-informative words of the source and target domains. The numbers in the Venn diagram indicate the number of words in both sets.
Phase 1 Data Accuracy
Source Target
Source Only 93.3 (0.1) 76.5 (0.2)
Target Only 92.9 (0.4) 82.3 (0.7)
Source + Target 93.5 (0.4) 83.3 (0.5)
Table 3: Comparison of CPT with varying data exposure during the first phase of training. Performance on the target domain strongly benefits more from exposure to target domain, and is boosted further with exposure to the source domain.
6 Why does Word Prediction Aid Classification for Generative UDA?
In this section, we examine why predicting masked words of the source and target domains through CPT boosts sentence classification on the unlabeled target domain for generative UDA. We hypothesize that by having to predict masked words that are informative to the downstream task during pre-training, the model implicitly learns information about the downstream task. For example, given the masked sentence, “I really _ the movie, it was a fascinating watch.”, the masked word is indicative of the downstream task, in this case sentiment analysis. The model can only predict this masked word (which would be a positive word like “loved” or “enjoyed”) by using other words informative to the task (“fascinating”). Through this process, the model is essentially learning features which are useful to the downstream task, despite having no direct supervision.
@@ 6 Why does Word Prediction Aid Classification for Generative UDA?
To test this hypothesis, we quantize the “informativeness” of each word to a classification task: an informative word is highly correlated with any of the labels in the downstream task. [^6] These informative words are similar to pivot features (Blitzer et al., 2006; Ziser & Reichart, 2018; Ben-David et al., 2020, inter alia), with the exception that they are chosen based on information from the source domain only. Specifically, we follow Gururangan et al. (2018) and use pointwise mutual information (PMI) (Fano, 1961) of the word with respect to the class label:
@@
where we count the frequency of a word-class pair on $\mathcal{D}_{\text{src}}$ to estimate $p(\text{word},\text{class})$, and similarly count a word and a class individually on $\mathcal{D}_{\text{src}}$ to estimate $p(\text{word})$ and $p(\text{class})$.
[…]
How Masking helps learn Classification
We first confirm that label-informative words indeed impact the classification performance of the CPT model. We do this by masking informative words from a sentence at inference. Figure 7 (a) shows us that the performance of the model on the source domain is not impacted by masking uninformative words, but drops on masking informative words. However, how do we know how much of this bias towards label informative words was learned during the continued pre-training phase, rather than during supervised fine-tuning? To attempt to disentangle the impact of the training phases, we train the model through selective masking (informative or uninformative) in the first phase of training, and minimize the impact of the second phase by making it a few-shot task. Figure 7 (c) shows us that the model performs best on classification when trained to predict label informative words during masking, indicating that the model does indeed learn features relevant to the downstream task during the first phase of training.
The interplay between CPT and Classification for Generative UDA
We now extend this analysis to the target domain to understand how CPT plays a role in learning features from the unlabeled domain. Figure 7 (d) shows us that informative masking outperforms uninformative masking by a significant gap, once again signaling that the masking process helps the model implicitly learn the downstream task. However, unlike with the source domain, random masking results in the strongest performance. This is due to the domain mismatch: the informative words for the source and target domains are not identical (Figure 6), and the supervised training on the source domain adds a bias towards source-informative words. The mixture of these two sets of words are best predicted through random masking, explaining its strong performance.
This phenomenon also draws the observation that random masking is preferred to selective masking for generative UDA, contrary to single domain settings where informative masking is more useful (Levine et al., 2021; Gu et al., 2020).
Refer to caption Refer to caption
Figure 7: The impact of selective masking on classification performance of a CPT trained model. Left: Masking label informative words during inference degrades classification performance, compared to the original unmasked sequence. However, the removal of uninformative words does not impact the model on the source domain. Right: The impact of continued pre-training can be partially disentangled from that of supervised training by making the supervised training phase few-shot and training the model to mask informative or uninformative words during the continued pre-training phase. Informative masking is most beneficial for the source domain, indicating that the model learns task relevant features during masking. On the target domain, informative masking still captures some knowledge about the downstream task, however, the supervised training phase adds a bias towards source label informative words. Thus, random masking is most powerful.
7 Discussion
The Computational Trade-off of CPT
Our results in Section 3.2 show that continued pre-training and methods promoting domain invariance are competitive with each other. Continued pre-training suffers from the computational drawback of requiring an additional phase of training. Conversely, invariance based methods are difficult to optimize, possibly requiring more runs to achieve a stable optimum, and having a higher amortized computational cost. Inspired by Karouzos et al. (2021), we introduce a simple single phase variant of continued pre-training which is equivalent in performance to its two phase variant, nullifying the additional computational overhead of the approach (more details in Appendix I).
UDA in the age of LLMs
Recent breakthroughs in scale have showcased that large language models (LLMs) are highly powerful, and can perform various downstream tasks with limited or no training. This may raise question on the relevance of the UDA problem as a whole — does a model even require expensive adaptation to a domain it may have already been exposed to during pre-training? In addition to our analysis in Section 5, we argue that the requirement to adapt small models to unseen domains still holds in specific cases. Small supervised models have been shown to outperform zero-shot general-purpose LLMs on various downstream tasks (Huang et al., 2023; Zhu et al., 2023), and safety critical domains like healthcare and finance would benefit more from these models than a generalist LLM. Our study does not address how to better adapt to domains, rather we investigate ways a model may adapt to data unseen during pre-training. This is a question that holds for current LLMs, and will continue to hold as long as models are unable to access infinite data during pre-training.
8 Related Work
UDA through Promoting Domain Invariance
A major class of approaches in Model-centric UDA methods (Ramponi & Plank, 2020) aims to minimize $\mathcal{H}\Delta\mathcal{H}$ divergence (Ben-David et al., 2010) between the source and target domain features, through adversarial training (Tzeng et al., 2014; Ganin et al., 2016; Tzeng et al., 2017; Guo et al., 2022, inter alia) or through minimizing measures of domain similarity (Bousmalis et al., 2016; Ge et al., 2023). Malik et al. (2023) have shown the minimization of MMD to outperform other invariance based methods. However, past work has shown that domain-invariance is a weak constraint for adaptation (Zhao et al., 2019; Karouzos et al., 2021), could introduce domain-specific hyperparameters (Trung et al., 2022), and is also prone to instability issues (Han & Eisenstein, 2019; Sun et al., 2019; Wilson & Cook, 2020; Kashyap et al., 2021).
UDA through Continued Pre-Training
The limitations of invariance-based model-centric methods have encouraged the emergence of alternate approaches, based on self-supervised learning through contrastive learning (Kumar et al., 2022; Shen et al., 2022; Long et al., 2022), pseudo-labeling (Zhou & Li, 2005; Ruder & Plank, 2017, inter alia) or language model pre-training. Despite not being directly useful to certain downstream tasks (Uppaal et al., 2023), CPT has been used for adaptation to labeled tasks, in both full fine-tuning (Gururangan et al., 2020; Lee et al., 2020; Gao et al., 2021, inter alia) and PEFT setups (Kim et al., 2021; Hung et al., 2023). A smaller body of work has explored the utility of CPT in a UDA setup (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Parović et al., 2023), identifying the class of methods to be more stable than invariance-based methods.
Generative UDA
The emergence of large language models (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia) introduced the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022; Gao et al., 2021; Liu et al., 2023), enabling multi-task training without any task specific architectural changes. However, the framework of casting discriminative classification tasks into generative next word prediction tasks has not yet been extended to UDA. The closest work to this setting (Ben-David et al., 2021) uses a generative model to create domain identifier prompts and feed them back into the model, however the final task label prediction is still discriminative. In our work, we focus on gaining insights to extend the powerful class of CPT methods to purely generative UDA, where prediction on the downstream task is treated as a next word prediction task. Through this, we also present novel findings on the impact of CPT to prompt-based classifiers in the UDA framework, countering previous findings from other studies in a single-domain setting (Gu et al., 2020; Levine et al., 2021; Wettig et al., 2023).